3 consecutive odd numbers add to give 405. Find the smallest of the numbers.
Answer by mananth(16949)

Let the numbers be x, x+2, x+4
their sum =405
..
x+x+2+x+4=405
3x+6=405
add -6
3x+6-6=405-6
3x=399
divide by 3
x= 133
the numbers are 133, 135,137.